OpenStudy (anonymous):

how do i show that two lines in 3D are the same?

OpenStudy (anonymous):

L1 : x = 1 + 3t, y = −2 + t, z = 2t L2 : x = 4 − 6t, y = −1 − 2t, z = 2 − 4t

OpenStudy (anonymous):

do i equate the 2 xs , the 2 ys, and the 2 zs like 1+3t=4-6t 9t=3 t=1/3 x=1+3(1/3)==>x=2 x=4-6(1/3)===>x=2 y=-2+(1/3)==>y=-5/3 y=-1-2(1/3)==>y=-5/3 and z=2(1/3)==>z=2/3 z=2-4(1/3)==>z=2/3 i think thats the way it shd be done....

OpenStudy (anonymous):

@welshfella

OpenStudy (welshfella):

looks good to me

OpenStudy (welshfella):

you have shown that the x , y and x values are same

OpenStudy (anonymous):

ok thanks